# Dual-Source SPIR over a noiseless MAC without Data Replication or Shared Randomness

### v1

@@ VI Concluding remarks
We studied information-theoretically secure SPIR in the absence of shared randomness, a noisy channel, and data replication. Instead, we leveraged a noiseless binary adder multiple-access channel and two non-colluding servers with independent content to provide information-theoretic security. We fully characterized the capacity region for this setting. While we assumed that all parties strictly follow the protocol, an open problem remains in addressing potential attempts by parties to cheat.
@@ Appendix A One-time pad lemma
=$\stackrel{{\scriptstyle(b)}}{{=}}I(A;B)+\epsilon+\delta,$ (30)
@@ Appendix B Proof of Lemma 1
$=I(Z_{1}Z_{2};\bar{C}_{1,1:K}\bar{C}_{2,1:K}|W_{1:K})+I(Y^{n}_{1:K}\mathbf{A}_{1:K};\bar{C}_{1,1:K}\bar{C}_{2,1:K}|Z_{1}Z_{2}W_{1:K})$
=$\stackrel{{\scriptstyle(a)}}{{=}}\sum_{t=1}^{K}I(Y^{n}_{t}\mathbf{A}_{t};\bar{C}_{1,1:K}\bar{C}_{2,1:K}|Z_{1}Z_{2}Y^{n}_{1:t-1}\mathbf{A}_{1:t-1}W_{1:K})$
$\stackrel{{\scriptstyle(b)}}{{\leqslant}}\sum_{t=1}^{K}H(Y^{n}_{t}\mathbf{A}_{t}|Z_{1}Z_{2}W_{t})-H(Y^{n}_{t}\mathbf{A}_{t}|\bar{C}_{1,1:K}\bar{C}_{2,1:K}Z_{1}Z_{2}Y^{n}_{1:t-1}\mathbf{A}_{1:t-1}W_{1:K})$
=$\stackrel{{\scriptstyle(c)}}{{=}}\sum_{t=1}^{K}H(Y^{n}_{t}\mathbf{A}_{t}|Z_{1}Z_{2}W_{t})-H(Y^{n}_{t}\mathbf{A}_{t}|\bar{C}_{1,t}\bar{C}_{2,t}Z_{1}Z_{2}W_{t})$
=$\stackrel{{\scriptstyle(d)}}{{=}}\sum_{t=1}^{K}H(Y^{n}_{t}\mathbf{A}_{t}|Z_{1}Z_{2}Z_{1,t}Z_{2,t}W_{t})-H(Y^{n}_{t}\mathbf{A}_{t}|\bar{C}_{1,t}\bar{C}_{2,t}Z_{1}Z_{2}Z_{1,t}Z_{2,t}W_{t})$
$\stackrel{{\scriptstyle(e)}}{{\leqslant}}\sum_{t=1}^{K}H(Y^{n}_{t}\mathbf{A}_{t}|Z_{1,t}Z_{2,t}W_{t})-H(Y^{n}_{t}\mathbf{A}_{t}|\bar{C}_{1,t}\bar{C}_{2,t}Z_{1,t}Z_{2,t}W_{t})$
@@
  6. and the limit holds by the Servers’ privacy with respect to the client in Line 7 of Algorithm 3.
@@ Appendix C Proof of Lemma 2
=$\stackrel{{\scriptstyle(a)}}{{=}}I(W_{1:K};F_{1,\mathcal{L}_{1}}F_{2,\mathcal{L}_{2}}|F_{1,Z_{1}}F_{2,Z_{2}}Z_{1,1:K}Z_{2,1:K})$
@@
$\stackrel{{\scriptstyle(b)}}{{\leqslant}}H((C_{1,t,i}[Z_{1,t,i}])_{t\in\llbracket 1,L_{1}-1\rrbracket\backslash\{Z_{1}\},i\in\llbracket 1,L_{2}-1\rrbracket}|F_{1,Z_{1}}F_{2,Z_{2}}Z_{1,1:K}Z_{2,1:K})$
@@
$\stackrel{{\scriptstyle(c)}}{{\leqslant}}(L_{1}-2)(L_{2}-1)\frac{nR_{1}}{L_{2}-1}+(L_{2}-2)(L_{1}-1)\frac{nR_{2}}{L_{1}-1}$
$\phantom{--}-H((S_{1,t,i})_{t\in\llbracket 1,L_{1}-2\rrbracket,i\in\llbracket 1,L_{2}-1\rrbracket}(S_{2,t,j})_{t\in\llbracket 1,L_{2}-2\rrbracket,j\in\llbracket 1,L_{1}-1\rrbracket})$
$=0,$
@@ Appendix D Proof of Lemma 3
$\stackrel{{\scriptstyle(a)}}{{\leqslant}}\sum_{t=1}^{K}H(\bar{C}_{1,t}X^{n}_{1,t}\mathbf{A}_{t})-H(\bar{C}_{1,t}X^{n}_{1,t}\mathbf{A}_{t}|Z_{1,t}Z_{2,t}\bar{C}_{1,1:t-1}X^{n}_{1,1:t-1}\mathbf{A}_{1:t-1})$
=$\stackrel{{\scriptstyle(b)}}{{=}}\sum_{t=1}^{K}H(\bar{C}_{1,t}X^{n}_{1,t}\mathbf{A}_{t})-H(\bar{C}_{1,t}X^{n}_{1,t}\mathbf{A}_{t}|Z_{1,t}Z_{2,t})$
@@
    holds by the Markov chain $(\bar{C}_{1,t},X^{n}_{1,t},\mathbf{A}_{t})-(Z_{1,t},Z_{2,t})-(\bar{C}_{1,1:t-1},X^{n}_{1,1:t-1},\mathbf{A}_{1:t-1})$, and the limit holds by the privacy of the client file selection.
Then, we have
@@
$=\sum_{t=1}^{K}I(\bar{C}_{1,t}X^{n}_{1,t}\mathbf{A}_{t};Z_{1,t}Z_{2,t}|\bar{C}_{1,1:t-1}X^{n}_{1,1:t-1}\mathbf{A}_{1:t-1})$
$\phantom{---}+I(\bar{C}_{1,t}X^{n}_{1,t}\mathbf{A}_{t};\bar{C}_{2,1:K}|Z_{1,t}Z_{2,t}\bar{C}_{1,1:t-1}X^{n}_{1,1:t-1}\mathbf{A}_{1:t-1})$
$\xrightarrow{n\to\infty}\sum_{t=1}^{K}I(\bar{C}_{1,t}X^{n}_{1,t}\mathbf{A}_{t};\bar{C}_{2,1:K}|Z_{1,t}Z_{2,t}\bar{C}_{1,1:t-1}X^{n}_{1,1:t-1}\mathbf{A}_{1:t-1})$
$=\sum_{t=1}^{K}[I(\bar{C}_{1,t}X^{n}_{1,t}\mathbf{A}_{t};\bar{C}_{2,t}|Z_{1,t}Z_{2,t}\bar{C}_{1,1:t-1}X^{n}_{1,1:t-1}\mathbf{A}_{1:t-1})$
$\phantom{---}+I(\bar{C}_{1,t}X^{n}_{1,t}\mathbf{A}_{t};\bar{C}_{2,1:K}|Z_{1,t}Z_{2,t}\bar{C}_{2,t}\bar{C}_{1,1:t-1}X^{n}_{1,1:t-1}\mathbf{A}_{1:t-1})]$
$\stackrel{{\scriptstyle(a)}}{{\leqslant}}\sum_{t=1}^{K}[I(\bar{C}_{1,t}X^{n}_{1,t}\mathbf{A}_{t};\bar{C}_{2,t}Z_{1,t}Z_{2,t})+I(\bar{C}_{1,t}X^{n}_{1,t}\mathbf{A}_{t};\bar{C}_{2,1:K}|Z_{1,t}Z_{2,t}\bar{C}_{2,t}\bar{C}_{1,1:t-1}X^{n}_{1,1:t-1}\mathbf{A}_{1:t-1})]$
@@
    $\leqslant H(\bar{C}_{1,t}X^{n}_{1,t}\mathbf{A}_{t})-H(\bar{C}_{1,t}X^{n}_{1,t}\mathbf{A}_{t}|Z_{1,t}Z_{2,t}\bar{C}_{2,t}\bar{C}_{1,1:t-1}X^{n}_{1,1:t-1}\mathbf{A}_{1:t-1})$

### v2

@@ VI Concluding remarks
We studied information-theoretically secure SPIR in the absence of shared randomness, a noisy channel, and data replication. Instead, we leveraged a noiseless binary adder multiple-access channel and two non-colluding servers with independent content to provide information-theoretic security. We fully characterized the capacity region for this setting. While we assumed that all parties strictly follow the protocol, an open problem is to address potential attempts by parties to cheat, for instance, as considered in [pinto2011achieving, dowsley2017oblivious, nascimento2008oblivious]. Other open problems include considering that the client has side information about the files stored in the servers, for instance, as considered in [chen2020capacity, 8882304, 8903451, 9216062, 9335998, 10061385, zivarifard2024private, zivarifard2025private, Shekofteh2025A], investigating achievable rates when only limited communication is allowed between the servers, and characterizing the capacity region for more than two servers. For the latter research direction, a first challenge is extending the converse, specifically generalizing Lemma 8 to more than two inputs. A second challenge arises in the achievability, in defining erasure sets that enable simultaneous private retrieval from all servers. Specifically, we conjecture that reducing the multi-file problem to a two-file-per-server scenario and repeatedly applying the two-file algorithm may not be optimal.
@@ Appendix A One-time pad lemma
$\stackrel{{\scriptstyle(b)}}{{\leqslant}}I(A;B)+\epsilon+\delta,$ (30)
@@ Appendix B Proof of Lemma 1
$=I(Z_{1}Z_{2};\bar{C}_{1,1:K}\bar{C}_{2,1:K}|W_{1:K})$
$\phantom{-}+I(Y^{n}_{1:K}\mathbf{A}_{1:K};\bar{C}_{1,1:K}\bar{C}_{2,1:K}|Z_{1}Z_{2}W_{1:K})$
$\stackrel{{\scriptstyle(a)}}{{=}}\sum_{t=1}^{K}I(Y^{n}_{t}\mathbf{A}_{t};\bar{C}_{1,1:K}\bar{C}_{2,1:K}|Z_{1}Z_{2}Y^{n}_{1:t-1}\mathbf{A}_{1:t-1}W_{1:K})$
$\stackrel{{\scriptstyle(b)}}{{\leqslant}}\sum_{t=1}^{K}H(Y^{n}_{t}\mathbf{A}_{t}|Z_{1}Z_{2}W_{t})$
$\phantom{--}-H(Y^{n}_{t}\mathbf{A}_{t}|\bar{C}_{1,1:K}\bar{C}_{2,1:K}Z_{1}Z_{2}Y^{n}_{1:t-1}\mathbf{A}_{1:t-1}W_{1:K})$
$\stackrel{{\scriptstyle(c)}}{{=}}\sum_{t=1}^{K}H(Y^{n}_{t}\mathbf{A}_{t}|Z_{1}Z_{2}W_{t})-H(Y^{n}_{t}\mathbf{A}_{t}|\bar{C}_{1,t}\bar{C}_{2,t}Z_{1}Z_{2}W_{t})$
$\stackrel{{\scriptstyle(d)}}{{=}}\sum_{t=1}^{K}H(Y^{n}_{t}\mathbf{A}_{t}|Z_{1}Z_{2}Z_{1,t}Z_{2,t}W_{t})$
$\phantom{--}-H(Y^{n}_{t}\mathbf{A}_{t}|\bar{C}_{1,t}\bar{C}_{2,t}Z_{1}Z_{2}Z_{1,t}Z_{2,t}W_{t})$
$\stackrel{{\scriptstyle(e)}}{{\leqslant}}\sum_{t=1}^{K}H(Y^{n}_{t}\mathbf{A}_{t}|Z_{1,t}Z_{2,t}W_{t})$
$\phantom{--}-H(Y^{n}_{t}\mathbf{A}_{t}|\bar{C}_{1,t}\bar{C}_{2,t}Z_{1,t}Z_{2,t}W_{t})$
@@
  6. (f)
    holds by the servers’ privacy with respect to the client in Line 7 of Algorithm 3.
@@ Appendix C Proof of Lemma 2
$\stackrel{{\scriptstyle(a)}}{{=}}I(W_{1:K};F_{1,\mathcal{L}_{1}}F_{2,\mathcal{L}_{2}}|F_{1,Z_{1}}F_{2,Z_{2}}Z_{1,1:K}Z_{2,1:K})$
@@
[…]
@@
$\stackrel{{\scriptstyle(c)}}{{\leqslant}}(L_{1}-2)(L_{2}-1)\frac{nR_{1}}{L_{2}-1}+(L_{2}-2)(L_{1}-1)\frac{nR_{2}}{L_{1}-1}-H((S_{1,t,i})_{t\in\llbracket 1,L_{1}-2\rrbracket,i\in\llbracket 1,L_{2}-1\rrbracket}(S_{2,t,j})_{t\in\llbracket 1,L_{2}-2\rrbracket,j\in\llbracket 1,L_{1}-1\rrbracket})$
$=0$ (31)
@@ Appendix D Proof of Lemma 3
$\stackrel{{\scriptstyle(a)}}{{\leqslant}}\sum_{t=1}^{K}H(\bar{C}_{1,t}X^{n}_{1,t}\mathbf{A}_{t})$
$\phantom{--}-H(\bar{C}_{1,t}X^{n}_{1,t}\mathbf{A}_{t}|Z_{1,t}Z_{2,t}\bar{C}_{1,1:t-1}X^{n}_{1,1:t-1}\mathbf{A}_{1:t-1})$
$\stackrel{{\scriptstyle(b)}}{{=}}\sum_{t=1}^{K}H(\bar{C}_{1,t}X^{n}_{1,t}\mathbf{A}_{t})-H(\bar{C}_{1,t}X^{n}_{1,t}\mathbf{A}_{t}|Z_{1,t}Z_{2,t})$
@@
    holds by the Markov chain $(\bar{C}_{1,t},X^{n}_{1,t},\mathbf{A}_{t})-(Z_{1,t},Z_{2,t})-(\bar{C}_{1,1:t-1},X^{n}_{1,1:t-1},\mathbf{A}_{1:t-1})$;
  3. (c)
    holds by the privacy of the client file selection.
Then, we have (33)
@@
$=\sum_{t=1}^{K}I(\bar{C}_{1,t}X^{n}_{1,t}\mathbf{A}_{t};Z_{1,t}Z_{2,t}|\bar{C}_{1,1:t-1}X^{n}_{1,1:t-1}\mathbf{A}_{1:t-1})+I(\bar{C}_{1,t}X^{n}_{1,t}\mathbf{A}_{t};\bar{C}_{2,1:K}|Z_{1,t}Z_{2,t}\bar{C}_{1,1:t-1}X^{n}_{1,1:t-1}\mathbf{A}_{1:t-1})$
$\stackrel{{\scriptstyle(a)}}{{=}}\sum_{t=1}^{K}I(\bar{C}_{1,t}X^{n}_{1,t}\mathbf{A}_{t};\bar{C}_{2,1:K}|Z_{1,t}Z_{2,t}\bar{C}_{1,1:t-1}X^{n}_{1,1:t-1}\mathbf{A}_{1:t-1})$
$=\sum_{t=1}^{K}[I(\bar{C}_{1,t}X^{n}_{1,t}\mathbf{A}_{t};\bar{C}_{2,t}|Z_{1,t}Z_{2,t}\bar{C}_{1,1:t-1}X^{n}_{1,1:t-1}\mathbf{A}_{1:t-1})+I(\bar{C}_{1,t}X^{n}_{1,t}\mathbf{A}_{t};\bar{C}_{2,1:K}|Z_{1,t}Z_{2,t}\bar{C}_{2,t}\bar{C}_{1,1:t-1}X^{n}_{1,1:t-1}\mathbf{A}_{1:t-1})]$
$\stackrel{{\scriptstyle(b)}}{{\leqslant}}\sum_{t=1}^{K}[I(\bar{C}_{1,t}X^{n}_{1,t}\mathbf{A}_{t};\bar{C}_{2,t}Z_{1,t}Z_{2,t})+I(\bar{C}_{1,t}X^{n}_{1,t}\mathbf{A}_{t};\bar{C}_{2,1:K}|Z_{1,t}Z_{2,t}\bar{C}_{2,t}\bar{C}_{1,1:t-1}X^{n}_{1,1:t-1}\mathbf{A}_{1:t-1})]$
@@
    $\leqslant H(\bar{C}_{1,t}X^{n}_{1,t}\mathbf{A}_{t})$
    $\phantom{-}-H(\bar{C}_{1,t}X^{n}_{1,t}\mathbf{A}_{t}|Z_{1,t}Z_{2,t}\bar{C}_{2,t}\bar{C}_{1,1:t-1}X^{n}_{1,1:t-1}\mathbf{A}_{1:t-1})$